Inverse laplace (s-5/s^2+4s+20)
1
Expert's answer
2021-07-16T15:07:14-0400

Solution;

Rewrite the denominator as follows;

s2+4s+20=s2+4s+4+16=(s+2)2+16

Rewrite the numerator as follows,

s-5=(s+2)-7

Therefore;

"L^{-1}[\frac{s-5}{s^2+4s+20}]" ="L^{-1}[\frac{s+2}{(s+2)^2+16}]" -"L^{-1}\frac{7}{(s+2)^2+16}"

="L^{-1}[\frac{s+2}{(s+2)^2+4^2}]-\frac74L^{-1}[\frac{4}{(s+2)^2+4^2}]"

From the Laplace tables;

f(t)=e-2tcos(4t)-"\frac 74"e-2tsin(4t)